Compton shift in wavelength is zero scattering angle of photon is:

a= 90

b=180

c=0

d=45


1
Expert's answer
2021-07-07T11:43:45-0400

We know that

"\lambda'-\lambda=\frac{h}{m_0c}(1-cos\phi)"

"\lambda=\frac{h}{m_0c}(1-cos\phi)""\rightarrow(1)"

Given

"∆\lambda=0"


equation (1) put value

"0=\frac{h}{m_0c}(1-cos\phi)"

"1-cos\phi=0"

"cos\phi=1"

"\phi=cos^{-1}(-1)=180°"

Option (b) is correct